# Geonhwa Jeong

### v2

@@ Abstract
Exploiting sparsity in deep neural networks (DNNs) has been a promising area to meet the growing computation need of modern DNNs. However, in practice, sparse DNN acceleration still faces a key challenge. To minimize the overhead of sparse acceleration, hardware designers have proposed structured sparse hardware support recently, which provides limited flexibility and requires extra model fine-tuning. Moreover, any sparse model fine-tuned for certain structured sparse hardware cannot be accelerated by other structured hardware.
To bridge the gap between sparse DNN models and hardware, this paper proposes tensor approximation via structured decomposition (TASD), which leverages the distributive property in linear algebra to turn any sparse tensor into a series of structured sparse tensors. Next, we develop a software framework, TASDER, to accelerate DNNs by searching layer-wise, high-quality structured decomposition for both weight and activation tensors so that they can be accelerated by any systems with structured sparse hardware support. Evaluation results show that, by exploiting prior structured sparse hardware baselines, our method can accelerate off-the-shelf dense and sparse DNNs without fine-tuning and improves energy-delay-product by up to 83% and 74% on average.
@@ 1 Introduction
DNNs have revolutionized various domains, such as computer vision [30, 38, 13], personal recommendation [44], speech recognition [4], and natural language processing [54, 12]. Meanwhile, DNN inference is also demanding more computation as they scale to billions of model parameters [15, 8, 60] and consume an enormous amount of input data [34, 9].
To address the increasing demand for DNN models, researchers propose to exploit sparsity in DNN models. Model pruning [19] is the most popular method to remove a set of parameters in the target DNN model based on certain criteria, inducing weight sparsity. This optimization exploits a phenomenon that empirically, large models are often overly parameterized and do not need all parameters to maintain the target accuracy at inference time. Deep learning (DL) model developers prefer to induce unstructured sparsity on weights so that they can focus on developing pruning algorithms to achieve better model accuracy and computation/model size trade-offs without any restriction. Also, many DNN models naturally exhibit activation sparisty due to the rectified linear unit (ReLU) that clips negative activation values to zeros, which induces unstructured activation sparsity.
@@
As a result, early sparse DNN accelerators [50, 64, 53] target unstructured sparsity and can accelerate any sparse DNNs. Unfortunately, unstructured sparsity support is challenging to deploy as a specialized HW unit as it causes significant area and energy overhead for indexing logic and flexible distribution/reduction logic [27]. Without native HW support, unstructured sparsity in DNNs lead to irregular memory accesses and diverged control/execution patterns, which are hostile to parallel architectures like GPUs [^1] The only sparse DNN accelerators on commercial devices is NVIDIA Sparse Tensor Core[46], which supports a fixed 2:4 structured sparsity. and specialized tensor accelerators like tensor core and systolic array. This forces DNN model developers today to constrain the DNN sparsity search space to only consider coarse-grained structured patterns, such as channel or filter pruning [42, 5], while maintaining iso-accuracy, thereby, moving the burden to DNN model developers.
To balance the need, recently, hardware (HW) designers [46, 27, 36, 72] have proposed fine-grained structured sparsity support. Such designs force DNN developers to induce sparsity with certain constraints at a cost of achievable sparsity degree for iso-accuracy comparison, but promise predictable performance gain on top of the existing DNN supports. For example, NVIDIA’s fine-grained 2:4 structured sparsity support balances the imposed constraints and achievable performance and provides a good target for DNN developers to optimize the network for. However, such methods impose an extra burden on DNN developers on top of the existing huge design space (DNN architectures, training recipes, etc.) to explore. Moreover, extending the existing support to more patterns [27, 62] means more overhead and design space exploration for both HW designers and DNN developers. We summarize different flows to exploit sparsity in Figure 1.
In this paper, we ask "Can we design a system to expose the flexible unstructured sparse interface to the DNN developers, but only with the efficient, less flexible structured sparse HW support?" Our answer is to introduce a new level of abstraction between DL model developers and hardware designers, similar to the abstraction layer in the instruction-set architectures. Specifically, we introduce a method, structured sparse tensor decomposition, to approximate any sparse tensor with a series of structured sparse tensors. Leveraging the distributed property of tensor algebra, we further propose to dynamically "decode" an unstructured sparse tensor algebra into a series of "microcode", i.e., structured sparse tensor algebra, which are efficient and compatible with prior structured sparse hardware.
@@
  • We present the first work to bridge unstructured sparse DNN and structured sparse HW with Tensor Approximation via Structured Decomposition (TASD), which approximate any sparse tensor with a series of structured sparse tensors.
  • We propose a framework, TASDER, which finds the TASD series for each DNN layer to accelerate dense/sparse DNNs with structured sparse hardware.
  • We propose a simple architectural extension and dataflow on top of existing structured sparse accelerators [27] to execute TASD series efficiently.
  • For various off-the-shelf dense and sparse DNNs, we show that TASD improves EDP by up to 83% and by 70% on average. We also show that across a range of DNNs, TASD can reduce the computation by 40%.
@@ 2.1 Terminology
Sparsity is a characteristic of data that includes zero values. The sparsity degree of a given tensor is measured as the fraction of the number of zeros in the tensor to the number of the total elements in the tensor. If a tensor has 0% sparsity degree, we call the tensor dense. Sparsity by itself is often used to indicate sparsity degree. To describe the zero distribution in a tensor, a sparsity pattern can be given to the tensor. If there is no defined sparsity pattern, we call it unstructured sparsity. Various patterns can be classified as structured sparsity, such as block sparse [43], butterfly sparse [10], and mixed patterns [69].
One of the most popular pattern is N:M structured sparsity [71], as it is supported in both commercial product [46] and academia proposals [27, 36] with active training recipe research [67]. An N:M structured sparse tensor means that there can be at most N non-zeros in each M-element block in a certain rank of the tensor as shown in Figure 2.
A view of a tensor $A$ for a sparsity pattern is a tensor after potentially dropping some elements to meet the rule of the sparsity pattern. For a matrix filled with non-zeros randomly, it is possible that the matrix does not meet the 2:4 sparsity pattern, i.e. there could be a block composed of 4 consecutive elements with more than two non-zeros. To generate a 2:4 view of the matrix, some non-zeros in the matrix should be dropped (pruned in DNNs) to meet the pattern. As this process could drop some original values, it can be lossy. Figure 2 also shows various tensors under different structured sparse views.
Prior work in DNN accelerators also proposed dense accelerators, unstructured sparse accelerators, 2:4 structured sparse accelerators, etc. To clarify the nomenclatures, in this paper, if a sparsity pattern is used to describe a hardware accelerator, such accelerator should provide the lossless and native support for any input tensor under such view.
@@ 2.2 DNN SW: Inducing sparsity in DNNs
Weight Sparsity. Without a specific pattern in mind, common model pruning methods introduces unstructured sparsity in weights [45]. Due to the irregular accesses to handle non-zeros in unstructured sparse matrices, unstructured sparsity is not adequate for accelerating DNNs on the existing parallel hardware, such as GPUs.
To address the issue, structured pruning forces pre-defined, static sparsity constraints in weights. For example, N:M structured sparsity [71, 14, 70, 25, 67, 62, 41] ensures that there are at most N non-zeros in each block composed of M consecutive elements, such that the required acceleration hardware can be trivial by exploiting the regularity in sparsity patterns. Thus, various accelerators, including recent sparse tensor cores in NVIDIA Ampere GPUs [46], target to exploit the fine-grained structured sparsity instead of unstructured sparsity. Nonetheless, structured pruning suffers from a higher loss of accuracy [16] than unstructured pruning since the extra pruning constraints reduce the flexibility. This extra loss of accuracy often leads to longer fine-tuning time (e.g., repeat the whole training process again) than unstructured sparse method to recover the loss in accuracy due to pruning [41].
Activation Sparsity. On the other hand, activation sparsity arises at runtime due to the non-linear activation functions such as ReLU, ReLU6 [22], and SquaredReLU [61], which clips negative values to zero. Activation sparsity is prevalent in both conventional Convolutional Neural Networks (CNNs) and recent Transformers [32]. Since it is intrinsic in DNN models, no extra fine-tuning or pruning steps are required to introduce activation sparsity. While the weight sparsity can be determined statically, activation sparsity is dynamic as the intermediate input feature map values depend on the inputs of the DNN model. Thus, the location of non-zeros and the degree of sparsity are unpredictable, similar to unstructured pruning, which makes it hard for structured sparse hardware to exploit input sparsity. Another challenge is that some recently proposed activation functions, such as GELU [21], and Swish [55], do not generate zero, which nullifies the benefits of exploiting activation sparsity in prior work [26].
@@ 2.3 DNN HW: Exploiting sparsity in DNNs
Table 1 shows the comparison of various sparse HW support for different DNN models. Prior unstructured sparse accelerators, including SCNN [50], SIGMA [53], Samsung NPU [26], and dual-side sparse core (DSTC) [64], target unstructured sparsity and skip redundant computations aggressively, but they suffer from non-trivial area/power costs due to the complex indexing and reduction logic, often introducing workload imbalance problems [37] as well. For example, SIGMA [53] introduces 37.7% area overhead compared to the dense baseline architecture due to its flexible and non-blocking distribution/reduction networks. Similarly, SCNN [50] and Griffin [59] produce 34% and 32% area overhead due to the support for the unstructured sparse dataflow. Moreover, when the sparsity degree is low or zero, they provide no improvement or even degrade performance and efficiency, due to the extra overhead for supporting unstructured sparsity.
@@
Dense [29, 47, 28]
Unstr [53, 50, 64]
@@
Table 1: Comparison of different DNN HW supports. Unstr: Unstructured sparse. Str: Structured sparse. Wgt: Weights. Act: Activations.
More recent structured sparse tensor accelerator architectures, such as STA [35], Sparse Tensor Core from NVIDIA GPUs (NV-STC) [46], and VEGETA [27], provide HW support for structured sparsity with minimal area overhead. However, these designs accelerate only structured pruned models that use the specific pattern supported. Also, they focus on weight sparsity since it is not trivial to exploit unstructured activation sparsity without much overhead. Recently, S2TA [37] has tried to circumvent the challenge to support sparse activation by forcing structured sparse pattern dynamically, but it requires modifying the existing models and even more fine-tuning steps.
@@ 2.4 Tension between sparse DNN SW and HW
Figure 3 shows the state of sparse DNN software and hardware. On the one hand, model developers have shown that unstructured sparsity provides a better model accuracy and a higher sparsity degree. On the other hand, hardware developers have shown that structured sparsity support is more practical to include in GPUs and other DNN accelerators. Such tension in the desired sparsity patterns hampers the progress in bringing sparse DNN acceleration to practice.
The main drawback of the previous hardware-specific patterns is that the pruning software and hardware support are tightly coupled, such that the software generates a model specifically pruned for the pattern supported by the hardware. For example, a model pruned for the NV-STC can only be accelerated by NV-STC, not by S2TA. To decouple the tight relationship, we propose another layer of system software between the model developers and DNN hardware for sparsity. Our insight is to approximate a tensor by decomposing it into a series of structured sparse tensors. We leverage the distributive property in tensor algebra to execute the series of structured sparse GEMM. This mechanism thus provides an unstructured sparse interface for developers but only requires structured sparse support from hardware. As shown in Table 1, by bridging DNN model and HW, this work is able to accelerate all types of sparsity seamlessly with a low area overhead.
Figure 3: TASD Interface.
@@ 3 TASD: Tensor Approximation via Structured Decomposition
We introduce a method to approximate unstructured sparsity using a series of structured sparsity. In this paper, we use a set of N:M structured sparsities for TASD to explain the method and show how to use it practically, but the concept is general and not limited to only N:M structured sparse patterns.
@@ 3.1 Overview
The extracted matrix, $A_{1}^{2:4}$ covers 70% in terms of the number of non-zero values while covering 84% in terms of the sum of the magnitudes. The percentage for the lost magnitudes is smaller than the percentage of the lost non-zero values because we extract two largest elements out of four consecutive elements. If we discard the remaining matrix $R_{1}$, then the original matrix $A$ can be approximated as $A_{1}^{2:4}$. Thus, we call this approximation structured decomposition. If we approximate $A$ with a 3:4 pattern instead of the 2:4 pattern, we can derive matrix $A_{1}^{3:4}$ with a structured decomposition that drops only one non-zero element, covering 90% of the number non-zeros and 96% of the sum of total magnitudes.
@@
We call $A$ as the original matrix and $\sum A_{i}^{s_{i}}$ as TASD series, to draw an analogy to the classic Taylor series [^2] Taylor series approximates any function with polynomials, while TASD series approximates any tensor with structured sparse tensors.: Each successive term (residual structure sparse tensor) improves the accuracy of the approximation. A TASD series configuration includes the number ("order") of TASD terms ($n$) and the structured sparsity pattern ($s_{i}$) for each TASD term. Using TASD, one can generate a structured sparse view of a given tensor, and the error between the view and the original tensor would depend on the TASD series configuration.
Using TASD for Matrix Multiplication: TASD decomposes any tensor $A$ into a series of structured sparse tensors. The decomposed tensors can be used in any tensor algebra, such as matrix multiplication ($C=A\times B$), which can be approximated as $A_{1}^{s_{1}}\times B$, so if $s_{1}$ is 2:4, and the matrix multiplication is running on NVIDIA STC, potentially 50% of Multiply-and-Accumulate (MAC) operations could be skipped.
With the distributive property of tensor algebra, matrix $A$ can be approximated using more TASD terms such as $(A_{1}^{s_{1}}+A_{2}^{s_{2}})B=A_{1}^{s_{1}}B+A_{2}^{s_{2}}B$. If $s_{1}$ is 2:4 and $s_{2}$ is 2:8, about 25% of MAC operations could be skipped. Thus, finding the proper TASD series to minimize the error while maximizing compute reduction will determine the quality of the approximation.
3.2 An analysis of TASD with synthetic data
We generate a synthetic matrix $B$ with dimensions of 128$\times$128 and densities ranging from 0.1 to 0.75. We explore three TASD series in this experiment; 1) using one term with 2:4, 2) using two terms with 2:4 and 2:8, and 3) using three terms with 2:4, 2:8, and 2:16. To consider various distributions, we tested two different distributions, a uniform random distribution between 0 and 1 and a normal distribution with a mean of 0 and a standard deviation of $\frac{1}{3}$. Figure 5 shows the results with matrices generated using the normal distribution.
Error Behavior: The first trend we observe is that the error gets smaller as the approximated sparsity gets lower since it is likely to drop fewer non-zeros with a more conservative approximation. Second, for the same approximated sparsity and the block size (M), the error gets smaller as the matrix A gets sparser. Given the same TASD configuration, the sparser matrices would drop fewer non-zeros using the same TASD configuration as shown in Figure 5, thus resulting in a smaller error. Third, with the same sparsity of matrix A and approximated sparsity, the N:4 configuration causes a larger error than the N:8 configuration (such as 1:4 and 2:8), since the expressiveness of the N:8 pattern is higher. Finally, given any unstructured sparse tensor, we can limit the error of matrix multiplication by conservatively selecting the TASD configuration, while maximizing the compute reduction. This optimization thus becomes the key to leveraging TASD for accelerating sparse DNN models with structured sparse hardware.

### v3

@@ Abstract
Exploiting sparsity in deep neural networks (DNNs) has been a promising area for meeting the growing computation requirements. To minimize the overhead of sparse acceleration, hardware designers have proposed structured sparsity support, but it provides limited flexibility and requires extra model fine-tuning. Moreover, any sparse model fine-tuned for certain structured sparse HW cannot be accelerated by other structured hardware. To enable acceleration using unstructured sparsity of DNNs on structured sparse hardware, we propose an approximation method leveraging the distributive property in linear algebra to turn any sparse tensor into a series of structured sparse tensors. We also develop a software framework, TASDER, to apply high-quality structured approximation on weights and activations of DNNs. Our method accelerates dense and sparse DNNs without fine-tuning and improves energy-delay-product (EDP) by up to 83% and 74%. It achieves up to 39% speed-up on a real system.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
DNNs have revolutionized various domains Krizhevsky et al. (2012); Liu et al. (2022a); Dosovitskiy et al. (2020); Naumov et al. (2019); Radford et al. (2019); Devlin et al. (2019), but DNN inference demands extreme compute power, memory footprints, and memory bandwidth as models scale to billions and trillions of parameters Shoeybi et al. (2019); Fedus et al. (2021); Chowdhery et al. (2022). To mitigate this growing demand, sparsity has emerged as a promising opportunity. Model pruning Han et al. (2015) is the most popular method to remove a set of parameters in DNNs. This optimization exploits a phenomenon that large models are often overly parameterized and do not need all parameters to maintain the target accuracy. Model pruning induces unstructured sparsity on weights (i.e. there is no specific pattern in the distribution of zeros) unless constraints are enforced during the pruning. Also, many DNN models naturally exhibit activation sparsity due to the rectified linear unit (ReLU) that clips negative activation values to zeros resulting in unstructured sparsity.
Unfortunately, unstructured sparsity in DNNs leads to irregular memory accesses and diverged control/execution patterns, which are hostile to parallel architectures like GPUs, tensor cores NVIDIA (2020b), and systolic arrays Jouppi et al. (2017). This has led to a plethora of academic research on unstructured sparsity support in DNN accelerators Parashar et al. (2017); Hegde et al. (2019); Qin et al. (2020); Wang et al. (2021); Gondimalla et al. (2023). Nonetheless, the complexity in hardware required for unstructured sparsity, such as indexing logic Hegde et al. (2019) and flexible distribution/reduction logic Qin et al. (2020), has been prohibitive in their commercial adoption.
Recent sparse DNN accelerators Zhu et al. (2019); NVIDIA (2020a); Liu et al. (2021); Jeong et al. (2023); AMD (2023); Xiao (2023) have proposed structured sparsity support. The most popular example is fine-grained N:M sparsity, which constrains at most N non-zeros in each block composed of M consecutive elements. A commercial example of this is NVIDIA’s sparse tensor core (STC) with 2:4 structured sparsity. Structured sparse hardware forces DNN model developers to induce sparsity with certain constraints. Otherwise, that hardware is unusable. For example, NVIDIA’s fine-grained 2:4 structured sparsity support could provide 2$\times$ throughput by skipping ineffectual computations only if the model is tuned with 2:4 structured sparsity. A model that does not follow 2:4 structured sparsity would, in fact, need to run on the dense tensor core, losing the benefits of sparsification. On the flip side, a 2:4 structured sparse model cannot be accelerated on different structured sparse hardware, such as a sparse tensor core with 1:4 or 2:8 structured sparsity support. Therefore, while the availability of structured sparse hardware is extremely promising, it remains limiting for model developers as it requires tuning the model to the specific sparsity support of the hardware, in addition to navigating the already vast design space, including DNN architectures and training recipes. This makes the broad adoption of structured sparse hardware challenging.
@@
In this paper, we ask and answer the following question “Can we decouple the dependency between DNN models and HW, so that a model with any sparsity pattern can be accelerated on any structured sparse HW?” Our answer is to introduce a new level of abstraction between DL model developers and hardware designers. Specifically, we introduce a method, structured sparse tensor decomposition, to approximate any sparse tensor (such as pruned weights or intrinsic sparse activations) with a series of structured sparse tensors. Leveraging the distributed property of tensor algebra, we further propose to dynamically “decode” an unstructured sparse tensor algebra into a series of structured sparse tensor algebra, which is efficient and compatible with prior structured sparse hardware. We contrast our proposed approach with current flows to exploit sparsity in Figure 1.
@@
  • We propose Tensor Approximation via Structured Decomposition (TASD), which approximates any sparse tensor with a series of structured sparse tensors. TASD is the first work to demonstrate running unstructured sparse DNNs on structured sparse HW.
  • We propose a framework, TASDER, a bridge between HW and SW to find the TASD series to accelerate dense/sparse DNNs on structured sparse hardware.
  • We propose a simple architectural extension and dataflow on top of existing structured sparse accelerators Jeong et al. (2023) to execute TASD series efficiently.
  • For various dense and sparse DNNs, we show that TASD improves EDP by up to 83% and by 70% on average. We also show that across a range of DNNs, TASD can reduce the computation by 40%.
  • We show the effectiveness of TASD on commercial hardware (NVIDIA RTX 3080 GPU) with 2:4 sparse tensor cores and achieve up to 39% performance gain.
@@ 2.1 Terminology
Sparsity is a characteristic of data that includes zeros. The sparsity degree of a given tensor is measured as the fraction of the number of zeros to the number of the total elements in the tensor. If a tensor has 0% sparsity degree, we call the tensor dense. Sparsity by itself is often used to indicate sparsity degree. To describe the distribution of zeros, a sparsity pattern can be given to the tensor. If there is no defined sparsity pattern, we call it unstructured sparsity. Various patterns can be classified as structured sparsity, such as block sparse Narang et al. (2017), butterfly sparse Dao et al. (2021), and mixed patterns Zaheer et al. (2020).
One of the most popular patterns is N:M structured sparsity Zhou et al. (2021), as it is supported in both commercial products NVIDIA (2020a); AMD (2023); Xiao (2023) and academic proposals Liu et al. (2021); Jeong et al. (2023); Bambhaniya et al. (2023) with active training recipe research Pool & Yu (2021); Mishra et al. (2021); NVIDIA (2021); Yazdanbakhsh et al. (2022); Lu et al. (2023); Frantar & Alistarh (2023); Bambhaniya et al. (2024). An N:M structured sparse tensor means that there can be at most N non-zeros in each M-element block in a certain rank of the tensor as shown in Figure 2 (c).
We define a view of a tensor $A$ for a sparsity pattern as a tensor after potentially dropping some elements to meet the rule of the sparsity pattern. For a matrix filled with non-zeros randomly, it is possible that the matrix does not meet the 2:4 sparsity pattern, i.e. there could be a block composed of 4 consecutive elements with more than two non-zeros. To generate a 2:4 view of the matrix, some non-zeros in the matrix should be dropped (pruned in DNNs) to meet the pattern. As this process could drop some original values, it can be lossy. Figure 2 also shows various tensors under different structured sparse views.
Prior work in DNN accelerators also proposed dense accelerators, unstructured sparse accelerators, 2:4 structured sparse accelerators, etc. To clarify the nomenclatures, in this paper, if a sparsity pattern is used to describe a hardware accelerator, that accelerator should provide lossless and native support for any input tensor under such view. For example, we call Google TPU a dense accelerator and NVIDIA Sparse Tensor Core a 2:4 accelerator.
@@ 2.2 DNN SW: Inducing sparsity in DNNs
Weight Sparsity. Without a specific pattern in mind, common model pruning methods introduces unstructured sparsity in weights Neuralmagic (2023). Due to the irregular accesses to handle non-zeros in unstructured sparse matrices, unstructured sparsity is not adequate for accelerating DNNs on the existing parallel hardware, such as GPUs.
To address the issue, structured pruning forces pre-defined sparsity constraints in weights. For example, N:M structured sparsity Zhou et al. (2021); Hubara et al. (2021); Sun et al. (2021); Mishra et al. (2021); Fang et al. (2022); Zhang et al. (2022); Yazdanbakhsh et al. (2022) ensures there are at most N non-zeros in each block composed of M consecutive elements, so the required acceleration hardware can be trivial by exploiting the regularity in sparsity patterns. Thus, various accelerators, including recent sparse tensor cores in NVIDIA Ampere GPUs NVIDIA (2020a), target to exploit the fine-grained structured sparsity instead of unstructured sparsity. Nonetheless, structured pruning suffers from a higher loss of accuracy Frantar & Alistarh (2023) than unstructured pruning since the extra pruning constraints reduce the flexibility. This extra loss of accuracy often leads to longer fine-tuning time (e.g., repeat the whole training process again) than unstructured sparse method to recover the loss in accuracy due to pruning Mishra et al. (2021).
Activation Sparsity. Activation sparsity arises at runtime due to the non-linear activation functions such as ReLU, ReLU6 Howard et al. (2017), and SquaredReLU So et al. (2021), which clips negative values to zero. Activation sparsity is prevalent in both conventional Convolutional Neural Networks (CNNs) and recent Transformers Li et al. (2023). Since it is intrinsic in DNN models, no extra fine-tuning or pruning steps is required to introduce activation sparsity. Unlike weight sparsity, activation sparsity is dynamic as the intermediate input activation values depend on the inputs of the DNN model. Thus, the location of non-zeros and the degree of sparsity are unpredictable, similar to unstructured pruning making it hard for structured sparse hardware to exploit input sparsity. Another challenge is that recently proposed activation functions, such as GELU Hendrycks & Gimpel (2016), and Swish Ramachandran et al. (2017), do not generate zero, which nullifies the benefits of exploiting activation sparsity in prior work Jang et al. (2021).
@@ 2.3 DNN HW: Exploiting sparsity in DNNs
Unstructured sparse accelerators, including SCNN Parashar et al. (2017), SIGMA Qin et al. (2020), Samsung NPU Jang et al. (2021), and dual-side sparse core (DSTC) Wang et al. (2021) target unstructured sparsity and skip redundant computations aggressively, but they suffer from non-trivial area/power costs due to the complex indexing and reduction logic, often introducing workload imbalance problems Liu et al. (2022b) as well. For example, SIGMA Qin et al. (2020) introduces 38% area overhead compared to the dense architecture due to its flexible and non-blocking distribution/reduction networks. SCNN Parashar et al. (2017) and Griffin Shin et al. (2021) produce 34% and 32% area overhead due to the support for the unstructured sparse dataflow. Moreover, when the sparsity degree is low or zero, they provide no improvement or even degrade performance/efficiency due to the overhead for supporting unstructured sparsity Wu et al. (2023). More recent structured sparse tensor accelerator architectures, such as STA Liu et al. (2020), Sparse Tensor Core from NVIDIA GPUs (NV-STC) NVIDIA (2020a), and VEGETA Jeong et al. (2023), provide HW support for structured sparsity with minimal area overhead. However, these designs accelerate only structured pruned models with the specific pattern and focus on weight sparsity since exploiting unstructured activation sparsity without much overhead is not trivial. S2TA Liu et al. (2022b) has tried to circumvent the challenge by forcing structured sparse patterns dynamically, but it requires modifying the models and even more fine-tuning steps. We compare different DNN HWs in Table 1.
Table 1: Comparison of DNN HWs. Unstr: Unstructured sparse. Str: Structured sparse. Wgt: Weights. Act: Activations.
@@
Dense Jouppi et al. (2017); NVIDIA (2020b); Jeong et al. (2021)
Unstr Qin et al. (2020); Parashar et al. (2017); Wang et al. (2021)
@@
Figure 3: TASD Interface.
@@ 2.4 Tension between sparse DNN SW and HW
Figure 3 shows the state of sparse DNN software and hardware. On the one hand, model developers have shown that unstructured sparsity provides better model accuracy and higher sparsity degree. On the other hand, hardware developers have shown that structured sparsity support is more practical to include in GPUs and other DNN accelerators. Such tension in the desired sparsity patterns hampers the progress in bringing sparse DNN acceleration to practice.
The main drawback of the previous hardware-specific patterns is that the pruning software and hardware support are tightly coupled, such that the software generates a model specifically pruned for the pattern supported by the hardware. For example, a model pruned for the NV-STC can only be accelerated by NV-STC, not by S2TA. To decouple the tight relationship, we propose another layer of system software between the model developers and DNN hardware for sparsity. Our insight is to approximate a tensor by decomposing it into a series of structured sparse tensors. We leverage the distributive property in tensor algebra to execute the series of structured sparse GEMM. This mechanism provides an unstructured sparse interface for developers but only requires structured sparse support from hardware. As shown in Table 1, by bridging DNN model and HW, this work is able to accelerate all types of sparsity seamlessly with a low area overhead.
@@ 3 TASD: Tensor Approximation via Structured Decomposition
In this section, we introduce a method to approximate unstructured sparsity using a series of structured sparsity, which we call TASD. In this paper, we use a set of N:M structured sparsities for TASD to explain the method and show how to use it practically, but the method is general and not limited to only N:M structured sparsity.
@@ 3.1 Overview
The extracted matrix, $A_{1}^{2:4}$ covers 70% in terms of the number of non-zero values while covering 84% in terms of the sum of the magnitudes. The percentage for the lost magnitudes is smaller than the percentage of the lost non-zero values because we extract two largest elements out of four consecutive elements. If we discard the remaining matrix $R_{1}$, then the original matrix $A$ can be approximated as $A_{1}^{2:4}$. Thus, we call this approximation, structured decomposition. If we approximate $A$ with a 3:4 pattern instead of the 2:4 pattern, we can derive matrix $A_{1}^{3:4}$ with a structured decomposition that drops only one non-zero element, covering 90% of the number non-zeros and 96% of the sum of total magnitudes.
@@
We call $A$ as the original matrix and $\sum A_{i}^{s_{i}}$ as TASD series, to draw an analogy to the classic Taylor series [^1] Taylor series approximates any function with polynomials, while TASD series approximates any tensor with structured sparse tensors.: Each successive term (residual structure sparse tensor) improves the accuracy of the approximation. A TASD series configuration includes the number (“order”) of TASD terms ($n$) and the structured sparsity pattern ($s_{i}$) for each TASD term. Using TASD, one can generate a structured sparse view of a given tensor, and the error between the view and the original tensor would depend on the TASD series configuration.
3.2 Using TASD for matrix multiplication
TASD decomposes any tensor $A$ into a series of structured sparse tensors. Decomposed tensors can be used in any tensor algebra, such as matrix multiplication ($C=A\times B$), which can be approximated as $A_{1}^{s_{1}}\times B$. If $s_{1}$ is 2:4, and the matrix multiplication is running on NVIDIA Sparse Tensor Cores, potentially 50% of Multiply-and-Accumulate (MAC) operations could be skipped.
With the distributive property of tensor algebra, matrix $A$ can be approximated using more TASD terms such as $(A_{1}^{s_{1}}+A_{2}^{s_{2}})B=A_{1}^{s_{1}}B+A_{2}^{s_{2}}B$. If $s_{1}$ is 2:4 and $s_{2}$ is 2:8, about 25% of MAC operations could be skipped. Finding the proper TASD series to minimize the error while maximizing compute reduction will determine the quality of the approximation. We provide a detailed analysis of TASD with synthetic data in Appendix A.